OpenStudy (anonymous):

Which of the following triangles is an obtuse triangle? Answer A. 6, 10, 12 B. 7, 13, 14 C. 8, 15, 17

OpenStudy (anonymous):

you can use Pythagorean Theorem on each triangle, test if each triangle will form a right triangle. If it does, its not an obtuse triangle.

OpenStudy (anonymous):

Example, for option A. \[\sqrt{6^{2}+10^{2}}=\sqrt{136}=11.66 or 12\]

OpenStudy (anonymous):

Thus, it forms a right triangle, try to do the same process for the other options. :))

OpenStudy (anonymous):

@sauravshakya is the answer B

OpenStudy (anonymous):

just check i under stand pythangan therom

OpenStudy (anonymous):

7, 13, 14

OpenStudy (anonymous):

so it is B

Parth (parthkohli):

The sides are listed in the form \(a,b,c\). The only work you have to do is to check if \(a^2 + b^2 < c^2\).

OpenStudy (anonymous):

Not quite. \[c^2 > a^2 + b^2\]\[14^2 > 13^2 + 7^2\]\[196 > 169 + 49\]\[196 \text{is not greater than } 218\]It's not B.

Parth (parthkohli):

If the statement holds true, then we have an obtuse triangle.

Parth (parthkohli):

Bad luck. Check if \(6^2 + 10^2 < 12^2\)

OpenStudy (anonymous):

No B is not the answer.

OpenStudy (anonymous):

i forgot to add D. 9,9,9√ 2

OpenStudy (anonymous):

If \(c^2 = a^2 + b^2\), then it's a right triangle. If \(c^2 > a^2 + b^2\), then it's an obtuse triangle. If \(c^2 < a^2 + b^2\), then it's an acute triangle. It's just a matter of testing the answer choices.

OpenStudy (anonymous):

oops. I made a mistake. Yes B is correct answer

OpenStudy (anonymous):

@sauravshakya It's not B. :P

OpenStudy (anonymous):

@Calcmathlete yeah but if you √ 218 its 14.76 and thats greater than 14

OpenStudy (anonymous):

THat's not the theorem though. What matters is if the SQUARED ones are bigger. Not the regular terms.

OpenStudy (anonymous):

Yeah B is AN ACUTE TRIANGLE

OpenStudy (anonymous):

yeah and its not a right triangle

Parth (parthkohli):

What are we even doing here? \( \color{Black}{\Rightarrow 7^2 + 13^2 < 14^2 }\) \( \color{Black}{\Rightarrow 49 + 169 < 196 }\) \( \color{Black}{\Rightarrow 218 < 196}\) Contradiction.

OpenStudy (anonymous):

Messi makes me nervous

OpenStudy (anonymous):

It's asking for an obtuse triangle. Not an acute triangle @sauravshakya

Parth (parthkohli):

It's just a repetition of what @Calcmathlete said.

OpenStudy (anonymous):

THE CORRECT ANSWER IS A.

OpenStudy (anonymous):

that make's it a right triangle

OpenStudy (anonymous):

Is the answer D.

OpenStudy (anonymous):

It is A like ^ said... \[c^2 > a^2 + b^2\]\[12^2 > 6^2 + 10^2\]\[144 > 36 + 100\]\[144 > 136\]See? The statement is true.

OpenStudy (anonymous):

so the answeer is a

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Yes...do you get why it is A though?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Alright. Just making sure...